OpenStudy (anonymous):

simplify the expression 10y − -2y − 3z − 9z − 6y

OpenStudy (anonymous):

combine like terms?

Directrix (directrix):

Yes, do that. 10y − -2y − 3z − 9z − 6y = 10y + 2y -3z -9z - 6y = 10y + 2y - 6y -3z - 9x = ?

OpenStudy (anonymous):

6y-3z-9x ?